OpenStudy (aa1996):

Graph the line with slope -2/3 passing through the point (-5,-4) . i forgot how to solve it i dont remmebr the formula . help please

OpenStudy (anonymous):

If you want to find the equation of line with a given a slope of which goes through the point (x1,y1), you can simply use the point-slope formula to find the equation: y-y1=m(x-x1) where m is the slope, and (x1,y1) is the point given. Plug in m=-2/3, x1=-5 and y1=-4 What is the equation you get?

OpenStudy (aa1996):

i got no equation but i think i remember it is y=mx+b

OpenStudy (anonymous):

No @AA1996 I'm asking you to plug in m=-2/3, x1=-5 and y1=-4 in this equation: y-y1=m(x-x1) ( this is the point-slope formula )

OpenStudy (aa1996):

you mean the formula is m=y1-y2/x1-x2.. somthing like that?

OpenStudy (anonymous):

No, Substitute m=-2/3, x1=-5 and y1=-4 in y-y1=m(x-x1) You get, y-(-4)=-2/3(x-(-5)) => y+4=-2/3(x+5) Simplify this to find the equation.

OpenStudy (aa1996):

oh ok i understand now thank u. sorry im kinda slow in those stuff

OpenStudy (anonymous):

Dont worry, it's okay:) Now, to draw a straight line, you have to find one more point, (point (-5,-4) is already given). You can find the point intercepting the x-axis, by equating x=0 in y+4=-2/3(x+5) @AA1996 Can you do it?

OpenStudy (aa1996):

yes i can I can do the rest now thanks

OpenStudy (anonymous):

@AA1996 oh good:)

OpenStudy (campbell_st):

if you know y = mx + b then the equation is y = -2/3 x + b substitute x = -5 and y = - 4 to find b

OpenStudy (anonymous):

Look @AA1996 here's a nice tutorial explaining how to graph a Point-Slope Form of a Line:)

OpenStudy (aa1996):

@kaushi THANK YOU SOOO MUCH!!! ur sooo much help

OpenStudy (anonymous):

@AA1996 You are welcome:)